Steve has 5 poker chips and 7 blue poker chips.Tom has 9 red poker chips and 3 blue poker chips. If steve's chips are worth a total of $69 and tom's are worth a total of $57, determinr the calue of a blue chip?
Answer by benni1013(165)

This is an application with the systems of equations.
Allow x to equal red chips and y to equal blue chips.
The set of equations will look like.
5x+7y=69
9x+3y=57
Multiply the top by 3 and the bottom by 7 to make the y terms the same.
15x+21y=207
63x+21y=399
Subtract and change the signs.
15x+21y=207
-63x-21y=-399
Add.
-48x=-192
Divide.
x=4
Substitute four into one of the equations.
5(4)+7y=69
20+7y=69
Subtract 20
7y=49
y=7
Since y is the blue chip the blue chip's value is $7.